OpenStudy (srslyjocelyn):

I've been stuck on this question for too long, I just don't understand it :/ 3x+y+4=9x

OpenStudy (michele_laino):

if we subtract \(9x\) from both sides, we get: \[3x + y + 4 - 9x = 9x - 9x\] and after a simplification we get: \[ - 6x + y + 4 = 0\] which is the equation of a line in the cartesian plane \((x,y)\)

OpenStudy (srslyjocelyn):

Do I then plug in numbers for x and y?

OpenStudy (michele_laino):

yes! If we substitute \(x=0\), then we get: \[ - 6 \cdot 0 + y + 4 = 0 \Rightarrow y = - 4\] so the line passes at point: \[{P_1} = \left( {0, - 4} \right)\] now, please substitute \(y=0\) nd then find \(x\)

OpenStudy (michele_laino):

and*

OpenStudy (mrnood):

However - since there are 2 unknowns and 1 equation you cannot 'solve' this with numbers the equation above is the best you can do, unless you are told more in the question

OpenStudy (srslyjocelyn):

So it would be -6x + 0 = 0 and you'd divide -6 on both sides so x would equal 0?

OpenStudy (srslyjocelyn):

The question's instructions are "Simplify the equation if possible. Then graph the equation by any appropriate method."

OpenStudy (michele_laino):

after the substitution, we can write: \[ - 6x + 0 + 4 = 0 \Rightarrow 6x = - 4\] @srslyjocelyn

OpenStudy (mrnood):

OK −6x+y+4=0 is the same as y = 6x -4 this is a straight line - you should be able to tell me the slope and intercept of the line and therefore graph it...

OpenStudy (michele_laino):

next I divide both sides by \(6\) so I can write this: \[\frac{{6x}}{6} = \frac{{ - 4}}{6}\] please simplify @srslyjocelyn

OpenStudy (srslyjocelyn):

x=-.6?

OpenStudy (michele_laino):

more precisely, we have: \[x = - \frac{2}{3}\] so the line passes at point \[{P_2} = \left( { - \frac{2}{3},0} \right)\]

OpenStudy (mrnood):

do you recognise y=mx +b as the form of a straight line?

OpenStudy (michele_laino):

if ytou want to draw such line, it is suffice to draw both points: \[{P_1} = \left( {0, - 4} \right),{P_2} = \left( { - \frac{2}{3},0} \right)\] and then an infinite line through such points

OpenStudy (srslyjocelyn):

Even more confused now wow... The issue is I honestly don't understand how I'd graph 2/3 or anything along those lines when the graph provided only allows whole number.

OpenStudy (michele_laino):

here is my drawing: